# Machine Unlearning: A Comprehensive Survey

### v2

@@ Abstract.
As the right to be forgotten has been legislated worldwide, numerous studies have sought to design unlearning methods to protect users’ privacy when they want to remove their data from machine learning service platforms. This has given rise to the concept of “machine unlearning” — a field dedicated to removing the influence of specified samples from trained models. This survey aims to systematically classify a wide range of machine unlearning studies, discussing their differences, connections, and open problems. We categorize current unlearning methods into four key areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in unlearning. Given that centralized unlearning is the primary focus, we further divide its discussion into two parts: first, at a higher level, we classify centralized unlearning into exact unlearning and approximate unlearning; second, we provide a detailed introduction to the techniques used in these methods. Next, we introduce federated unlearning, another emerging area that has garnered significant attention. Following the discussion on unlearning methods, we review studies on unlearning verification and audit, which assess the effectiveness of existing unlearning algorithms. Additionally, we consider the privacy and security issues essential in machine unlearning and compile the related literature. Finally, we discuss the challenges of various unlearning scenarios and highlight potential research directions.
Federated Unlearning; Machine Unlearning; Unlearning Verification
ccs: Computing methodolgies Machine learningccs: Computing methodolgies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
@@ 1. Introduction
Machine unlearning has drawn growing research attention as the recent legislation of the ”Right to be Forgotten” in many countries. Notable instances include the GDPR in the European Union (Mantelero, 2013), the PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, ), and the California Consumer Privacy Act in the United States (de la Torre, 2018). According to these laws, companies must take reasonable measures to guarantee that personal data is deleted upon request. It indicates that individual users have the right to request companies to remove their private data, which was previously collected for ML model training. The deletion is not only erasing their data from a database, but it also needs to delete the influence of the specified samples from trained models. The process of data removal from models was first conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, suppose a user (Alice) wants to exercise her right (Mantelero, 2013) when quitting a ML application, then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the dataset previously used for ML model training is requested to be deleted; second, the ML model provider erases the contribution of these data from the trained models. A naive data-erasing method is retraining a new ML model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
@@
Many researchers have tried to find efficient and effective methods to implement unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. (1) Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training results non-deterministic (Bourtoule et al., 2021) and raises challenges for machine unlearning to estimate the impact of the typical erased samples. (2) Incrementality of training: The training process in machine learning is incremental, meaning that the model update from one data point influences the contribution of subsequent data points fed into the model. Deciding a way to effectively remove the contributions of the to-be-erased samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). (3) Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model typically has worse model utility than the model retrained from scratch. The degradation would be exponential, especially when a method tries to delete a huge amount of data samples. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies mitigate model utility degradation by bounding the loss function or restricting the unlearning update threshold, eliminating catastrophic unlearning remains an open problem. Recently, many studies have put efforts into solving these three main challenges and proposed many novel mechanisms that promote the progress of machine unlearning.
[…]
@@
Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Graph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024; Wu et al., 2023b)Federated (Distributed) Unlearning: (Wang et al., 2022a, 2024; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
@@
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify the current unlearning literature into four main scenarios: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences and connections, as well as their advantages and disadvantages. We first categorize unlearning approaches into four main areas: centralized unlearning, federated unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig. 1. Since many studies focus on centralized unlearning, we dedicate two sections to introducing corresponding studies in this area. We classify these unlearning methods into: Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost. Approximate unlearning tries to unlearn a model approximate to the model retrained using the remaining dataset. We also provide a detailed introduction to the popular techniques used to implement these two types of unlearning. In addition to centralized unlearning methods, we introduce federated (distributed) unlearning studies (Wang et al., 2022a; Liu et al., 2021), which are gaining increasing importance. After discussing unlearning implementation methods, verifying the unlearning effect is equally crucial and has garnered significant research attention (Thudi et al., 2022b). Therefore, we subsequently review related research on unlearning verification. Finally, we consider privacy and security issues in machine unlearning essential, organizing and reviewing the related publications on privacy and security threats, defenses, and unlearning applications.
@@
There are only a few surveys on machine unlearning because it is a relatively new research domain, and most of them are preprints on Arxiv, focusing on certain unlearning scenarios. For an introduction to machine unlearning, including discussions on exact and approximate unlearning problems and their solutions through recently proposed methods, see (Mercuri et al., 2022). For information on provable machine unlearning for linear models, including algorithm introductions and experimental analysis, refer to (Mahadevan and Mathioudakis, 2021). For an overview of federated unlearning, see (Liu et al., 2023b; Wang et al., 2023d), and for graph unlearning, see (Said et al., 2023). While Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it, they focused primarily on introducing problem formulations and technical definitions. However, none of these surveys reviewed the privacy and security issues in machine unlearning, nor did they comprehensively discuss the differences and connections between various unlearning scenarios.
@@
Compared with existing surveys on machine unlearning, we performed a systematic literature review (SLR) following the guidelines presented by (Keele et al., 2007) and as exemplified in (Kubrak et al., 2022). The contributions of this paper include:
@@
  • We systematically catalog machine unlearning studies according to their scenarios, which can be briefly classified into centralized unlearning, distributed unlearning, unlearning verification, and privacy and security issues in machine unlearning.
@@
  • For the centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
@@
  • We discuss different challenges between various machine unlearning areas, from centralized unlearning to federated unlearning, unlearning verification, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
The survey’s remaining sections are arranged as follows. We show a systematic literature review model of the survey in Section 2. The basic knowledge of machine unlearning and related technical tools are summarized in Section 3. After introducing the background, we introduce the survey following the taxonomy order in Fig. 1. The primary content of centralized unlearning is presented in Section 4, which includes an introduction from two main unlearning categories and corresponding in-depth techniques. We collect and introduce the federated unlearning scenarios in Section 5. Section 6 introduces the unlearning evaluation and verification methods. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy and security threats accompanied by machine unlearning. Section 8 discusses machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way. After that, we enumerate the unaddressed issues for further research. At last, in Section 10, we provide a summary of the survey.
@@ 3.1. Machine Unlearning
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2, and we summarize the notations in Table 1. Let $\mathcal{Z}$ denote as a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
The unlearning process begins at an unlearning request when a user wants to erase his specified data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples (Bourtoule et al., 2021), classes (Warnecke et al., 2023), or graph nodes (Chen et al., 2022b). Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$. The unlearned model can be described as $M_{D\setminus D_{e}}=\mathcal{U}(M,D,D_{e})$. When dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effect, as in Step 3 in Fig. 2. Therefore, we add the verification step as the last step in the unlearning process, and we will introduce the related literature later.
Figure 2. Machine Unlearning Process

### v3

@@ Abstract.
As the right to be forgotten has been legislated worldwide, numerous studies have sought to design unlearning methods to protect users’ privacy when they want to remove their data from machine learning service platforms. This has given rise to the concept of “machine unlearning” — a field dedicated to removing the influence of specified samples from trained models. This survey aims to systematically classify a wide range of machine unlearning studies, discussing their differences, connections, and open problems. We categorize current unlearning studies into four key areas: traditional unlearning methods, unlearning verification, domain-centric machine unlearning, and privacy and security issues in unlearning. In the traditional unlearning part, we first classify traditional unlearning into exact unlearning and approximate unlearning at a higher level; then, we provide detailed introductions to the techniques used in these studies. Next, we review studies on unlearning verification to introduce how to assess the effectiveness of unlearning operations. Following the discussion of traditional unlearning and verification methods, we introduce domain-centric machine unlearning, including graph unlearning, federated unlearning, diffusion model unlearning, and large language model unlearning. Additionally, we consider privacy and security issues essential in machine unlearning and compile related literature. Finally, we discuss the challenges of various unlearning scenarios and highlight potential research directions.
Machine Unlearning; Unlearning Verification; Federated Unlearning; Graph Unlearning; Diffusion Model Unlearning; Large Language Model Unlearning
$c$s: Computing methodologies Machine learningccs: Computing methodologies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
@@ 1. Introduction
Machine unlearning has drawn growing research attention as the recent legislation of the ”Right to be Forgotten” in many countries. Notable instances include the GDPR in the European Union (Mantelero, 2013), the PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, ), and the California Consumer Privacy Act in the United States (de la Torre, 2018). According to these laws, companies must take reasonable measures to guarantee that personal data is deleted upon request. It indicates that individual users have the right to request companies to remove their private data, which was previously collected for ML model training. The deletion is not only erasing their data from a database, but it also needs to delete the influence of the specified samples from trained models. The process of data removal from models was conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, suppose a user (Alice) wants to exercise her right (Mantelero, 2013) when quitting a ML application, then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the dataset previously used for ML model training is requested to be deleted; second, the ML model provider erases the contribution of these data from the trained models. A naive data-erasing method is retraining a new ML model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
@@
Many researchers have tried to find efficient and effective methods to implement unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. (1) Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training results non-deterministic (Bourtoule et al., 2021) and raises challenges for machine unlearning to estimate the impact of the typical erased samples. (2) Incrementality of training: The training process in machine learning is incremental, meaning that the model update from one data point influences the contribution of subsequent data points fed into the model. Deciding a way to effectively remove the contributions of the to-be-erased samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). (3) Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model typically has worse model utility than the model retrained from scratch. The degradation would be severe, especially when a method tries to delete a huge amount of data samples. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies mitigate model utility degradation by bounding the loss function or restricting the unlearning update threshold, eliminating catastrophic unlearning remains an open problem. Recently, many studies have put efforts into solving these three main challenges and proposed many novel mechanisms that promote the progress of machine unlearning.
@@
Machine UnlearningTraditional Unlearning MethodsExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a; Schelter et al., 2023; Koch and Soll, 2023; Hu et al., 2024a)Split Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022a)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021; Tarun et al., 2023; Wang et al., 2023a; Chundawat et al., 2023; Warnecke et al., 2024; Liu et al., 2023a; Lin et al., 2023; Cha et al., 2024) Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2023)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Unlearning Verification: (Hu et al., 2022; Guo et al., 2023; Thudi et al., 2022a; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a; Thudi et al., 2022b) Domain-Centric Machine UnlearningGraph Unlearning: (Chien et al., 2022a; Chen et al., 2022b; Li et al., 2024a; Cong and Mahdavi, 2022; Wang et al., 2023b; Wu et al., 2023a; Zhang et al., 2024a; Wu et al., 2023b)Federated (Distributed) Unlearning Methods: (Wang et al., 2022a, 2024a; Fraboni et al., 2024; Wang et al., 2023c; Liu et al., 2021, 2022a; Zhang et al., 2023a; Gupta et al., 2021; Wu et al., 2022b)Diffusion Model Unlearning: (Alberti et al., 2025; Chen et al., 2024; Wu et al., 2025a; Li et al., 2024b; Zhang et al., 2024b; Wu et al., 2025b; Cywiński and Deja, 2025; Zhang et al., 2024c; Ko et al., 2024; George et al., 2025)Large Language Models Unlearning: (Yao et al., 2024a, b; Liu et al., 2025a, 2024a; Huang et al., 2024a; Wang et al., 2025a; Liu et al., 2024b; Jin et al., 2024; Li et al., 2024c)Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b; Hu et al., 2024b; Zhang et al., 2023b)Security Threats in Unlearning : (Marchant et al., 2022; Zhao et al., 2024; Di et al., 2022; Hu et al., ; Liu et al., 2024c)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Wei et al., 2023; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
@@
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify the current unlearning literature into four main scenarios: traditional unlearning methods, unlearning verification, domain-centric machine unlearning, and privacy and security issues in machine unlearning.
@@
This survey aims to classify and systematize machine unlearning methods based on the research problems and objectives in the unlearning process and to review their differences, connections, as well as their advantages and disadvantages. Our survey includes four main categories: traditional unlearning methods, unlearning verification, domain-centric unlearning methods, and privacy and security issues in machine unlearning, as shown in Fig. 1. In the traditional unlearning category, we will introduce the representative Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022a; Nguyen et al., 2020) methods, including basic concepts and detailed technical implementation. After discussing traditional unlearning methods, verifying the unlearning effectiveness is equally crucial and has garnered significant research attention (Thudi et al., 2022b). Therefore, we subsequently review related research on unlearning verification. Next, we will introduce domain-centric unlearning methods, including graph unlearning (Chien et al., 2022a; Chen et al., 2022b), federated unlearning (Wang et al., 2022a; Liu et al., 2021), diffusion model unlearning (Alberti et al., 2025; Chen et al., 2024), and large language model (LLM) unlearning (Yao et al., 2024a, b). Finally, we consider privacy and security issues in machine unlearning essential, organizing and reviewing the related publications on privacy and security threats, defenses, and unlearning applications.
Table 1. Comparison of this survey with representative existing surveys on machine unlearning.
Survey Year Main focus Difference from this survey
Mahadevan and Mathioudakis, Certifiable Machine Unlearning for Linear Models (Mahadevan and Mathioudakis, 2021) 2021 Experimental review of approximate/certifiable unlearning methods for linear models, with emphasis on efficiency, effectiveness, and certifiability trade-offs. We provide a broader taxonomy spanning traditional unlearning, verification, domain-centric unlearning, such as federated unlearning and LLM unlearning, and privacy/security issues.
Nguyen et al., A Survey of Machine Unlearning (Nguyen et al., 2022b) 2022 Broad overview of formulations, design criteria, removal requests, algorithms, scenarios, and applications. Offers a strong general foundation, but places more emphasis on concepts and formulations than on systematically comparing scenario-specific challenges.
Wang et al., Federated Unlearning and Its Privacy Threats (Wang et al., 2023d) 2023 Federated unlearning, especially unlearner roles, privacy leakage, membership inference risks, and defenses. Focused on federated settings only, particularly privacy threats in federated unlearning, rather than the full machine unlearning landscape.
Liu et al., A Survey on Federated Unlearning: Challenges, Methods, and Future Directions (Liu et al., 2023b) 2024 Dedicated taxonomy of federated unlearning methods, workflows, challenges, applications, and future directions. Comprehensive for federated unlearning, but not designed to compare federated unlearning with centralized unlearning, verification, and broader privacy/security issues within one unified framework.
Liu et al., Threats, Attacks, and Defenses in Machine Unlearning: A Survey (Liu et al., 2025b) 2025 Security-oriented review of vulnerabilities, attacks, and defenses in machine unlearning systems. Centers on threat taxonomy and defensive mechanisms, rather than on the full algorithmic and scenario-based landscape of machine unlearning.
This survey Unified, SLR-based survey covering traditional centralized unlearning, unlearning verification, domain-centric unlearning, privacy/security issues, applications, open questions, and their interconnections. Distinctive in integrating previously separate strands into one taxonomy and explicitly comparing their differences, connections, advantages, limitations, and research gaps.
@@
There are only a few surveys on machine unlearning because it is a relatively new research domain, and most of them are preprints on Arxiv, focusing on certain unlearning scenarios. For an introduction to machine unlearning, including discussions on exact and approximate unlearning problems and their solutions through recently proposed methods, see (Mercuri et al., 2022). For information on provable machine unlearning for linear models, including algorithm introductions and experimental analysis, refer to (Mahadevan and Mathioudakis, 2021). For an overview of federated unlearning, see (Liu et al., 2023b; Wang et al., 2023d), and for graph unlearning, see (Said et al., 2023). While Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it, they focused primarily on introducing problem formulations and technical definitions. For an security-oriented review of vulnerabilities in machine unlearning systems, see (Liu et al., 2025b). We provide a comparison difference between existing surveys in Table 1.
@@
Compared with existing surveys on machine unlearning, we performed a systematic literature review (SLR) following the guidelines presented by (Keele et al., 2007) and as exemplified in (Kubrak et al., 2022). The main contributions are as follows.
@@
  • We systematically catalog machine unlearning studies into traditional unlearning methods, unlearning verification, domain-centric unlearning, and privacy and security issues in machine unlearning.
@@
  • For the traditional centralized unlearning scenario, which draws the most attention, we divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
@@
  • We discuss different challenges in various machine unlearning areas, from traditional centralized unlearning to federated unlearning and generative model unlearning, unlearning verification, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
The survey’s remaining sections are arranged as follows. We show a systematic literature review model of the survey in Section 2. The basic knowledge of machine unlearning and related technical tools are summarized in Section 3. After introducing the background, we introduce the survey following the taxonomy order in Fig. 1. The primary content of traditional unlearning is presented in Section 4, which includes two main unlearning categories and corresponding in-depth techniques. Section 5 introduces the unlearning evaluation and verification methods. We collect and introduce the unlearning methods for specific domains such as federated unlearning, graph unlearning, diffusion model unlearning, and LLMs unlearning, in Section 6.1. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy and security threats accompanied by machine unlearning. Section 8 discusses machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way and enumerate the potential directions for further research. At last, in Section 10, we provide a summary of the survey.
@@ 3.1. Machine Unlearning
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2, and we summarize the notations in Table 2. Let $\mathcal{Z}$ denote a space of data items; the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
The unlearning process begins at an unlearning request when a user wants to erase his specified data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples (Bourtoule et al., 2021), classes (Warnecke et al., 2023), or graph nodes (Chen et al., 2022b). Then, in Step 2, the server removes the contribution of $D_{e}$ using a designed machine unlearning algorithm $\mathcal{U}$. The unlearned model can be described as $M_{D\setminus D_{e}}=\mathcal{U}(M,D,D_{e})$. The standard aim of unlearning is to ensure the unlearned model $\mathcal{U}(M,D,D_{e})$ is the same as the retrained model $\mathcal{A}(D\setminus D_{e})$ (i.e., $\mathcal{U}(M,D,D_{e})\simeq\mathcal{A}(D\setminus D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effectiveness, as in Step 3 in Fig. 2. Therefore, we add the verification step as the last step in the unlearning process, and we will introduce the related literature later.
@@
Figure 2. Overview of the machine unlearning pipeline, from model training and unlearning request types to unlearning execution and verification methods.